# Towards Context-aware Reasoning-enhanced Generative Searching in E-commerce

### v2

@@ Abstract.
[…]
@@ 1. Introduction
Search-based recommendation is one of the most critical application scenarios of modern e-commerce platforms, enabling users to discover and compare products among millions of offerings efficiently. The search results directly influence purchase decisions and platform revenue, making its effectiveness crucial for both user satisfaction and business success. Consequently, improving the quality of search-based recommendations has emerged as a key driver of user engagement, conversion rates, and long-term loyalty in online retail ecosystems.
@@
Contextual information constitutes an essential component of search-based recommendation, as it captures users’ implicit preferences that complement explicit query terms, especially when query expressions are noisy or incomplete, which makes it difficult to infer the true underlying intent from the surface form alone. Rich contextual cues—such as sequential dependencies between adjacent searches (e.g., a user searching for a smartphone followed by a phone case) or the emphasis on particular attributes—can reveal latent connections among target items. Moreover, user profiles, geographic locations, and other behavioral signals provide valuable priors about users’ interest distributions, akin to the collaborative filtering principle of leveraging similar behaviors across user groups. Consequently, effectively extracting and modeling such complex contexts during the search process remains a significant challenge for building more accurate and personalized search-based recommendation systems.
@@
Numerous efforts have been devoted to improving search performance from both matching-based and generative perspectives. The former typically focuses on enhancing query–item relevance through semantic retrieval or embedding models. Representative works include classical lexical-matching methods such as TF-IDF and BM25 (Salton and Buckley, 1988; Robertson et al., 2009). With the advent of deep learning, subsequent research has proposed neural architectures and tailored training strategies to learn discriminative embeddings for queries and targets (Huang et al., 2013; Karpukhin et al., 2020; Khattab and Zaharia, 2020). More recently, general-purpose embedding models pretrained on large-scale corpora have been explored (Neelakantan et al., 2022; Xiao et al., 2024), demonstrating superior performance across diverse retrieval scenarios. In contrast, generative approaches reformulate the search task as a sequence generation problem. Early studies transform items into semantic IDs and model the correlation between a query and its target in an autoregressive manner (Tay et al., 2022; Rajput et al., 2023), with subsequent improvements achieved through diverse aspects like multi-modal information fusion (Luo et al., 2024; Yang et al., 2025) and explicit injection of ranking ability (Li et al., 2024). Despite these advances, existing solutions still struggle to achieve a comprehensive understanding of complex search contexts and therefore to better infer user intentions, which remains a major obstacle to building more accurate and personalized search-based recommendation systems.
@@
To address these challenges, we propose a context-aware, reasoning-enhanced generative searching framework, namely CRS, which explicitly leverages the world knowledge and multi-step reasoning capabilities of large language models (LLMs) to deeply understand complex e-commerce search contexts. Concretely, we unify heterogeneous user and item contextual signals into a compact text-based representation or text-semantic identifiers and align them, which allows LLMs to operate directly on a common semantic substrate. Operating on this unified representation, the LLM performs explicit reasoning about latent user intent and its associations with item attributes, recommending candidate items in a generative, end-to-end fashion. To instill stable and effective reasoning behaviors without requiring dense human supervision, we develop a self-evolving post-training paradigm that iteratively alternates supervised fine-tuning (SFT) with reinforcement learning (RL), allowing the model to bootstrap and refine its own reasoning trajectories. Finally, recognizing that standard RL formulations can introduce ranking bias in search scenarios, we introduce a debiased variant of GRPO (Shao et al., 2024) tailored to generative searching to improve ranking fidelity and training stability. Together, these components enable explicit, context-aware reasoning that closes the gap between noisy queries and users’ true intentions, yielding a single end-to-end model that better exploits rich contextual signals for search-based recommendation. Our major contributions are as follows:
@@ 2.1. Matching-based Search
Deep semantic matching has gradually replaced purely term-based retrieval in modern e-commerce and information retrieval. Traditional lexical matching methods, including TF–IDF weighting and the BM25 ranking function (Salton and Buckley, 1988; Robertson et al., 2009), compute relevance scores by comparing exact or near-exact term overlaps between queries and documents and have long served as the backbone of large-scale search engines. Building on these foundations, early neural approaches such as the Deep Structured Semantic Model (DSSM) and its subsequent variants (Huang et al., 2013; Yu et al., 2021) employ dual-tower architectures that encode queries and items into a shared embedding space, thereby enabling efficient similarity computation and large-scale retrieval beyond surface term matching. More recently, pre-trained language models have been directly adapted to retrieval (Zhao et al., 2024). DPR (Karpukhin et al., 2020) learns dense query and document representations and consistently outperforms sparse baselines in open-domain QA and general search. Hybrid or late-interaction models, such as ColBERT and SPLADE (Khattab and Zaharia, 2020; Formal et al., 2021), further enhance fine-grained token-level matching while retaining scalable retrieval properties. Moreover, recent works explore building general embedding models (Neelakantan et al., 2022; Xiao et al., 2024; Zhang et al., 2025a) pre-trained on a large corpus of sentence pairs with contrastive learning, yielding promising retrieval performance in various domains. Despite these advances, the vast majority of matching-based methods still treat each query largely in isolation and devote limited attention to modeling rich contextual factors, which constrains their ability to capture users’ latent interests and dynamic search intentions.
$2.$ Generative Recommendation and Search
@@
The generative modeling paradigm has also been extensively explored in the search domain. GENRE (De Cao et al., 2020) first demonstrates the feasibility of using a pretrained T5 model to autoregressively generate target entity titles for retrieval, thereby casting entity search as a sequence generation problem. This line of work is further advanced by DSI and NCI (Tay et al., 2022; Wang et al., 2022), which introduce a learn-to-encode objective and a prefix-aware weight-adaptive decoder to improve retrieval accuracy and scalability. In parallel, several studies integrate learning-to-rank principles into generative retrieval to strengthen ranking quality. For example, LTRGR (Li et al., 2024) continues training a pretrained generative retriever with a margin-based loss, whereas ListGR (Tang et al., 2024) incorporates supervision derived from the Plackett–Luce model into the generation process. In the e-commerce setting, RARE (Liu et al., 2025a) introduces the notion of commercial intents to guide generation, and GRAM (Pang et al., 2025) designs 16 predefined textual identifier codes to represent items, thereby achieving better alignment with commercial text and background knowledge. More recently, the OneSearch (Chen et al., 2025) model has been proposed as the first industrially deployed end-to-end generative framework for large-scale e-commerce search, achieving significant improvements.
Moreover, with the rapid improvement of the reasoning capabilities of large language models (LLMs) (Guo et al., 2025), an emerging line of work has begun to investigate reasoning-enhanced recommendation from both latent and explicit reasoning perspectives. LARES (Liu et al., 2025b) and LatentR3 (Zhang et al., 2025b) improve representation capacity by performing additional computations in a latent reasoning space without relying on high-quality chain-of-thought data. By contrast, R2ec (You et al., 2025) unifies generative reasoning and discriminative recommendation in a single framework guided purely by reinforcement learning.
@@ 3.1.1. Representation
Furthermore, to allow the framework to generate target items without heavy post-processing, we convert item contexts into compact semantic IDs (SID) (Tay et al., 2022; Rajput et al., 2023). Specifically, we first transform the JSON texts into initial item embeddings $E^{0}=\{e_{1}^{0},e_{2}^{0},\cdots,e_{\mathcal{I}}^{0}\}$ using the bge-small-zh-v1.5 model (Xiao et al., 2024), where $\mathcal{I}$ denotes the full item set. We then apply a residual K-Means quantization algorithm (Deng et al., 2025; Luo et al., 2024), which ensures the coverage of codebook’s each layer by design, to obtain multi-layer SIDs. In detail, the generation process of layer $l$ is expressed as:
@@
Here $C^{l}$ is the codebook of the $l$-th layer and NN denotes the codebook size. In practice, we employ a codebook of $K=3$ layers augmented with an additional deduplication layer. The codes selected from each layer are concatenated to yield a unique four-token SID for every item (e.g., <a_23><b_1><c_124><d_0>). These SIDs of four layers are further registered as new special tokens in the LLM’s vocabulary, thereby fully integrating item identifiers into the unified textual representation.
@@ 3.2. Self-evolving Post-training
Concretely, we initialize from an SFT model $\mathcal{M}_{sft}^{0}$ obtained by further tuning the aligned model $\mathcal{M}_{align}$ on a small set of context analysis traces distilled from Qwen3-30B-A3B-Instruct-2507 (Team, 2025). This preliminary distillation step mainly enhances the base model’s instruction-following and slow-thinking ability. We then sample a subset of training instances to efficiently fine-tune the model with RL, thereby encouraging exploration. The details of RL training are explained in Section $3.$ Inspired by the filtering strategy of DAPO (Yu et al., 2025), we run the model on the training corpus and collect only those cases it gets wrong; training on these “hard” examples incentivizes the model to discover effective policies precisely where its contextual understanding is weak. After obtaining an RL-updated model $\mathcal{M}_{rl}^{0}$, we perform a subsequent SFT stage to explicitly consolidate the acquired knowledge, i.e., exploitation, by training on the correctly solved instances along with their generated reasoning trajectories. This yields a refined SFT model $\mathcal{M}_{sft}^{1}$, thus completing one iteration of the self-evolving loop. The training procedure consists of several iterations.
@@
$3.$ Debiased Reinforcement Learning
@@ 3.3.2. Debiased Algorithm: R-GRPO
Concretely, we decouple the traditional rollout into two stages: (i) generation of reasoning trajectories and (ii) beam search of predicted items. Given a context $q$, the model first generates $G$ context-aware reasoning trajectories $\{o_{i}\}_{i=1}^{G}$. For each trajectory, we then conduct an independent beam search to obtain a ranked list of $K$ candidate items, resulting in a total of $G\cdot K$ rollouts $\{o_{i}^{j}\}_{i=1,j=1}^{G,K}$. Each rollout is assigned a scalar reward $\{r_{i}^{j}\}_{i=1,j=1}^{G,K}$ according to our reward system. To more effectively integrate ranking information into RL training, for each reasoning trajectory we select the item prediction with the highest reward as the optimal outcome, and compute a rank-aware, weighted-sum reward to reflect the entire ranked list. That is,
@@
Note that we focusing on modifying the rollout process and the corresponding reward estimation schema to inject the ranking information, therefore our approach can be easily extended to diverse variants of recent RL algorithms (Yu et al., 2025; Zheng et al., 2025).
@@ 4.1.1. Datasets
Though there exist some open-source search-based recommendation datasets (Sun et al., 2023; Liu et al., 2023), they lack the essential non-anonymous contexts required by our approach. Therefore to evaluate the effectiveness of the proposed CRS framework, we curated three datasets from the log data of a large-scale e-commerce platform. Each dataset contains rich and complex search contexts, including user profile features, spatial-temporal information, historical interactions (including the click and non-click interactions), and metadata of the current query. Specifically, based on the user’s target item, we randomly sampled 100K and 50K users across all product categories, and additionally collected 27K users from the fashion category. For every user, we retained the 10 most recent interactions to capture short-term behavioral context. This process yields three benchmark datasets, denoted as All-100K, All-50K, and Fashion-27K, respectively. The detailed statistics of these datasets are summarized as follows in Table 1.
@@
4.1.2. Baselines
To comprehensively assess the effectiveness of the proposed CRS framework under the search-based recommendation scenario, we compare it against a diverse set of baselines covering both traditional match-based retrieval and recent generative search models:
  • BGE (Xiao et al., 2024) and Qwen3-Embedding (Zhang et al., 2025a) are unified embedding models pretrained on large-scale corpora and used to retrieve candidate items via embedding similarity. Here we use the bge-small-zh-v1.5 and Qwen3-Embedding-0.6B model for embedding computing.
  • LatentR3 (Zhang et al., 2025b) adopts a latent reasoning paradigm to further improve generative recommendation quality.
4.1.3. Experiments Details and Evaluation Protocols
To ensure a fair comparison among all generative models, we adopt a unified item codebook configuration. The number of codebook layers $K$ is fixed to 3, with layer sizes of 512, 256, and 256 for the three datasets, respectively. For the LLM backbones, we use mT5-base for GENRE, DSI, and LTRGR, and Qwen3-0.6B for LatentR3 and our CRS framework, following the settings recommended in their original papers. In addition, similar to CRS, LatentR3 takes user context as input, whereas the other approaches rely solely on the current query text. It is also worth noting that comparable performance is observed for GENRE, DSI, and LTRGR when their backbones are replaced with Qwen3-0.6B.
@@ 4.1.3. Experiments Details and Evaluation Protocols
Figure 2. Performance evolution of the proposed CRS framework during self-evolving post-training on the All-100K and All-50K datasets, measured by NDCG@10 and HR@10.
For evaluation, we split the datasets by user to generate the training and evaluation sets. The same prefix trie of our framework is utilized to guide the beam search process of all the generative models to guarantee the fairness of evaluation. We employ two widely used top-NN metrics: NDCG@NN (Normalized Discounted Cumulative Gain) and HR@NN (Hit Rate). NN is set to 1, 5, 10 for a comprehensive evaluation on both retrieval and ranking performance. All models are implemented in PyTorch and optimized using the Adam optimizer.

### v3

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792562isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Retrieval models and rankingccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Search-based recommendation is one of the most critical application scenarios of modern e-commerce platforms, enabling users to discover and compare products among millions of offerings efficiently (Chen et al., 2025; Cheng et al., 2022; Mao et al., 2024). The search results directly influence purchase decisions and platform revenue, making its effectiveness crucial for both user satisfaction and business success. Consequently, improving the quality of search-based recommendations has emerged as a key driver of user engagement, conversion rates, and long-term loyalty in online retail ecosystems.
@@
Contextual information constitutes an essential component of search-based recommendation, as it captures users’ implicit preferences that complement explicit query terms, especially when query expressions are noisy or incomplete, which makes it difficult to infer the true underlying intent (Cheng et al., 2026). Rich contextual cues—such as sequential dependencies between adjacent searches (e.g., a user searching for a smartphone followed by a phone case) or the emphasis on particular attributes—can reveal latent connections among target items. Moreover, user profiles, geographic locations, and other behavioral signals provide valuable priors about users’ interest distributions, akin to the collaborative filtering principle of leveraging similar behaviors across user groups. Consequently, effectively extracting and modeling such complex contexts during the search process remains a significant challenge for building more accurate and personalized search-based recommendation systems.
@@
Numerous efforts have been devoted to improving search performance from both matching-based and generative perspectives. The former typically focuses on enhancing query–item relevance through semantic retrieval or embedding models. Representative works include classical lexical-matching methods such as TF-IDF and BM25 (Salton and Buckley, 1988; Robertson et al., 2009). With the advent of deep learning, subsequent research has proposed neural architectures and tailored training strategies to learn discriminative embeddings for queries and targets (Huang et al., 2013; Karpukhin et al., 2020; Khattab and Zaharia, 2020). More recently, general-purpose embedding models pretrained on large-scale corpora have been explored (Neelakantan et al., 2022; Xiao et al., 2024), demonstrating superior performance across diverse retrieval scenarios. In contrast, generative approaches reformulate the search task as a sequence generation problem. Early studies transform items into semantic IDs and model the correlation between a query and its target in an autoregressive manner (Tay et al., 2022; Rajput et al., 2023), with subsequent improvements achieved through diverse aspects like multi-modal information fusion (Luo et al., 2024; Yang et al., 2025) and explicit injection of ranking ability (Li et al., 2024). Despite these advances, existing solutions still struggle to achieve a comprehensive understanding of complex search contexts, which remains a major obstacle to building more accurate and personalized search-based recommendation systems.
@@
To address these challenges, we propose a context-aware, reasoning-enhanced generative searching framework, namely CRS, which explicitly leverages the world knowledge and multi-step reasoning capabilities of large language models (LLMs) (Luo et al., 2025; Cheng et al., 2025; Guo et al., 2025a), to deeply understand complex search contexts. Concretely, we unify heterogeneous user and item contextual signals into a compact text-based representation or text-semantic identifiers and align them, which allows LLMs to operate directly on a common semantic substrate. Operating on this unified representation, the LLM performs explicit reasoning about latent user intent and its associations with item attributes, recommending candidate items in a generative, end-to-end fashion. To instill stable and effective reasoning behaviors without requiring dense human supervision, we develop a self-evolving post-training paradigm that iteratively alternates supervised fine-tuning (SFT) with reinforcement learning (RL), allowing the model to bootstrap and refine its own reasoning trajectories. Finally, recognizing that standard RL formulations can introduce ranking bias in search scenarios, we introduce a debiased variant of GRPO (Shao et al., 2024) tailored to generative searching to improve ranking fidelity and training stability. Together, these components enable explicit, context-aware reasoning that closes the gap between noisy queries and users’ true intentions, yielding a single end-to-end model that better exploits rich contextual signals for search-based recommendation. Our major contributions are as follows:
@@ 2.1. Matching-based Search
Deep semantic matching has gradually replaced purely term-based retrieval in modern e-commerce and information retrieval. Traditional lexical matching methods, including TF–IDF weighting and the BM25 ranking function (Salton and Buckley, 1988; Robertson et al., 2009), compute relevance scores by comparing exact or near-exact term overlaps between queries and documents and have long served as the backbone of large-scale search engines. Building on these foundations, early neural approaches such as the Deep Structured Semantic Model (DSSM) and its subsequent variants (Huang et al., 2013; Yu et al., 2021) employ dual-tower architectures that encode queries and items into a shared embedding space, thereby enabling efficient similarity computation and large-scale retrieval beyond surface term matching. More recently, pre-trained language models have been directly adapted to retrieval (Zhao et al., 2024). DPR (Karpukhin et al., 2020) learns dense query and document representations and consistently outperforms sparse baselines in open-domain QA and general search. Hybrid or late-interaction models, such as ColBERT and SPLADE (Khattab and Zaharia, 2020; Formal et al., 2021), further enhance fine-grained token-level matching while retaining scalable retrieval properties. Moreover, recent works explore building general embedding models (Neelakantan et al., 2022; Xiao et al., 2024; Zhang et al., 2025b) pre-trained on a large corpus of sentence pairs with contrastive learning, yielding promising retrieval performance in various domains. Despite these advances, the vast majority of matching-based methods still treat each query largely in isolation and devote limited attention to modeling rich contextual factors, which constrains their ability to capture users’ latent interests and dynamic search intentions.
2.2. Generative Recommendation and Search
@@
The generative modeling paradigm has also been extensively explored in the search domain. GENRE (De Cao et al., 2020) first demonstrates the feasibility of using a pretrained T5 model to autoregressively generate target entity titles for retrieval, thereby casting entity search as a sequence generation problem. This line of work is further advanced by DSI and NCI (Tay et al., 2022; Wang et al., 2022), which introduce a learn-to-encode objective and a prefix-aware weight-adaptive decoder to improve retrieval accuracy and scalability. In parallel, several studies integrate learning-to-rank principles into generative retrieval to strengthen ranking quality. For example, LTRGR (Li et al., 2024) continues training a pretrained generative retriever with a margin-based loss, whereas ListGR (Tang et al., 2024) incorporates supervision derived from the Plackett–Luce model into the generation process. In the e-commerce setting, RARE (Liu et al., 2025b) introduces the notion of commercial intents to guide generation, and GRAM (Pang et al., 2025) designs 16 predefined textual identifier codes to represent items, thereby achieving better alignment with commercial text and background knowledge. More recently, the OneSug (Guo et al., 2025b) and OneSearch (Chen et al., 2025) are the first models proposed as the first industrially deployed end-to-end generative framework for large-scale e-commerce search, achieving significant improvements.
Moreover, with the rapid improvement of the reasoning capabilities of large language models (LLMs) (Guo et al., 2025a), an emerging line of work has begun to investigate reasoning-enhanced recommendation from both latent and explicit reasoning perspectives. LARES (Liu et al., 2025a) and LatentR3 (Zhang et al., 2025a) improve representation capacity by performing additional computations in a latent reasoning space without relying on high-quality chain-of-thought data. By contrast, R2ec (You et al., 2025) unifies generative reasoning and discriminative recommendation in a single framework guided purely by reinforcement learning.
@@ 3.1.1. Representation
Furthermore, to allow the framework to generate target items without heavy post-processing, we convert item contexts into compact semantic IDs (SID) (Tay et al., 2022; Rajput et al., 2023). Specifically, we first transform the JSON texts into initial item embeddings $E^{0}=\{e_{1}^{0},e_{2}^{0},\cdots,e_{\mathcal{I}}^{0}\}$ using the bge-small-zh-v1.5 model (Xiao et al., 2024), where $\mathcal{I}$ denotes the full item set. We then apply a residual K-Means quantization algorithm (Deng et al., 2025; Luo et al., 2024) to obtain multi-layer SIDs. Compared to the widely used RQ-VAE approach, this design naturally ensures full codebook coverage without training, avoiding the collapse issue and making it suitable for large-scale SID construction. In detail, the generation process of layer $l$ is expressed as:
@@
Here $C^{l}$ is the codebook of the $l$-th layer and $N$ denotes the codebook size. In practice, we employ a codebook of $K=3$ layers augmented with an additional deduplication layer. The codes selected from each layer are concatenated to yield a unique four-token SID for every item (e.g., <a_23><b_1><c_124><d_0>). These SIDs of four layers are further registered as new special tokens in the LLM’s vocabulary, thereby fully integrating item identifiers into the unified textual representation.
@@ 3.2. Self-evolving Post-training
Concretely, we initialize from an SFT model $\mathcal{M}_{sft}^{0}$ obtained by further tuning the aligned model $\mathcal{M}_{align}$ on a small set of context analysis traces distilled from Qwen3-30B-A3B-Instruct-2507 (Team, 2025). This preliminary distillation step mainly enhances the base model’s instruction-following and slow-thinking ability. We then sample a subset of training instances to efficiently fine-tune the model with RL, thereby encouraging exploration. The details of RL training are explained in Section 3.3. Inspired by the filtering strategy of DAPO (Yu et al., 2025), we run the model on the training corpus and collect only those cases it gets wrong; training on these “hard” examples incentivizes the model to discover effective policies precisely where its contextual understanding is weak. After obtaining an RL-updated model $\mathcal{M}_{rl}^{0}$, we perform a subsequent SFT stage to explicitly consolidate the acquired knowledge, i.e., exploitation, by training on the correctly solved instances along with their generated reasoning trajectories. This yields a refined SFT model $\mathcal{M}_{sft}^{1}$, thus completing one iteration of the self-evolving loop. The training procedure consists of several iterations.
@@
3.3. Debiased Reinforcement Learning
@@ 3.3.2. Debiased Algorithm: R-GRPO
Concretely, we decouple the traditional rollout into two stages: (i) generation of reasoning trajectories and (ii) beam search of predicted items. Given a context $q$, the model first generates $G$ context-aware reasoning trajectories $\{o_{i}\}_{i=1}^{G}$. For each trajectory, we then conduct an independent beam search to obtain a ranked list of $K$ candidate items, resulting in a total of $G\cdot K$ rollouts $\{o_{i}^{j}\}_{i=1,j=1}^{G,K}$. Each rollout is assigned a scalar reward $\{r_{i}^{j}\}_{i=1,j=1}^{G,K}$ according to our reward system. To more effectively integrate ranking information into RL training, for each reasoning trajectory, we select the item prediction with the highest reward as the optimal outcome and compute a rank-aware, weighted-sum reward to reflect the entire ranked list. That is,
@@
Note that we are focusing on modifying the rollout process and the corresponding reward estimation schema to inject the ranking information, therefore our approach can be easily extended to diverse variants of recent RL algorithms (Yu et al., 2025; Zheng et al., 2025). We summarize the complete training workflow of R-GRPO in Section A to better clarify the procedure of the training process.
@@ 4.1.1. Datasets
Though there exist some open-source search-based recommendation datasets (Sun et al., 2023; Liu et al., 2023), they lack the essential non-anonymous contexts required by our approach. Therefore, to evaluate the effectiveness of the proposed CRS framework, we curated three datasets from the log data of Kuaishou’s e-commerce platform. Each dataset contains rich and complex search contexts, including user profile features, spatial-temporal information, historical interactions (including the click and non-click interactions), and metadata of the current query. Specifically, based on the user’s target item, we randomly sampled 100K and 50K users across all product categories, and additionally collected 27K users from the fashion category. For every user, we retained the 10 most recent interactions to capture short-term behavioral context. This process yields three benchmark datasets, denoted as All-100K, All-50K, and Fashion-27K, respectively. The detailed statistics of these datasets are summarized as follows in Table 1.
@@
4.1.2. Baselines
To comprehensively assess the effectiveness of the proposed CRS framework in search-based recommendation, we compare it against a diverse set of baselines covering both traditional match-based retrieval and recent generative search models:
  • BGE (Xiao et al., 2024) and Qwen3-Embedding (Zhang et al., 2025b) are unified embedding models pretrained on large-scale corpora and used to retrieve candidate items via embedding similarity. Here we use the bge-small-zh-v1.5 and Qwen3-Embedding-0.6B model for embedding computing.
  • LatentR3 (Zhang et al., 2025a) adopts a latent reasoning paradigm to further improve generative recommendation quality.
4.1.3. Experiments Details and Evaluation Protocols
To ensure a fair comparison across all generative models, we adopt a unified item codebook configuration. Specifically, the number of codebook layers $K$ is fixed to 3, with layer sizes set to 512, 256, and 256 for the three datasets, respectively. For the LLM backbones, we employ mT5-base for GENRE, DSI, and LTRGR, and Qwen3-0.6B for LatentR3 and our proposed CRS framework, following the configurations recommended in their original implementations. All models are trained until convergence over multiple epochs. In addition, consistent with CRS, LatentR3 incorporates user context as input, whereas the remaining methods rely solely on the current query text. Notably, we observe comparable performance for GENRE, DSI, and LTRGR when their backbones are replaced with Qwen3-0.6B.
For evaluation, we split the datasets by user to generate the training and evaluation sets. The same prefix trie of our framework is utilized to guide the beam search process of all the generative models to guarantee the fairness of evaluation. We employ two widely used top-$N$ metrics: NDCG@$N$ (Normalized Discounted Cumulative Gain) and HR@$N$ (Hit Rate). $N$ is set to 1, 5, 10 for a comprehensive evaluation on both retrieval and ranking performance. All models are implemented in PyTorch and optimized using the Adam optimizer.